OpenStudy (anonymous):

What is the value of the function y=2x-3 when x = –1? A. –5 B. 2 C. –1 D. 3

OpenStudy (anonymous):

is it A -5?

OpenStudy (anonymous):

What is the value of the function v=3/2t-5/2 when t = 3? is it 2

OpenStudy (anonymous):

Evaluate the function y = 15 – x when x = –4 A. –60 B. 19 C. 11 D. –60 is it B

OpenStudy (anonymous):

Yes,it is !

OpenStudy (anonymous):

Find the value of the function when n = 8. p = 0.009n A. 0.0072 B. 0.72 C. 0.072 D. 8.009 Is it C

OpenStudy (anonymous):

What is the value of the function when x = 4? y = 3x + 7 is it 19?

OpenStudy (anonymous):

am i right on all of them

OpenStudy (anonymous):

\(\bf\huge\color{#ff0090}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) @germansiser15

OpenStudy (anonymous):

thanks am i right on all of them though

OpenStudy (anonymous):

yes, it is !

OpenStudy (anonymous):

ok thank you sooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo much